A model rocket is launched with an acceleration of 22.5 m/s2 at an angle of 72.0o (measured from the horizontal). If the model rocket has a mass of 3.55 kg what would be the net horizontal and vertical forces involved
1
Expert's answer
2020-10-23T11:51:08-0400


Let the acceleration magnitude be "a = 22.5m\/s^2" and the mass "m = 3.55kg". Then, according to the Newton's second law, the net force acting on the model is colinear with the acceleration and has the magnitude (see the picture):

"F = ma = 3.55\times22.5 = 79.875\space N"

The horizontal component of this force is:


"F_x = F\cos\theta = 79.875\times \cos72\degree \approx 24.68\space N"

The vertical component:


"F_y = F\sin\theta = 79.875\times \sin72\degree \approx 75.97\space N"

Answer. Horizontal 24.68 N, vertical 75.97 N.